2. An object, initially at rest, starts to fall freely. What will be the speed of the object 10 seconds later? Find acceleration of the object in 5 seconds.
Answer #1

Initial speed of the object, u = 0 m/s (starts from rest)

time, t = 10 seconds

acceleration of free falling object, a = 9.8 m/s2

v = u + at = 9.8*10 = 98 m/s

The acceleration of the object after 5 seconds is equal to g = 9.8 m/s2